OpenStudy (anonymous):

Use algebra to find the inverse of the given one-to-one function. f(x)=5+(square root of 2x -5)

OpenStudy (anonymous):

\[f(x)=5+\sqrt{2x-5}\]

OpenStudy (anonymous):

x=((y-5)^2+5)/2

OpenStudy (anonymous):

did you get it?

OpenStudy (anonymous):

thats wronf according to my hw site

OpenStudy (anonymous):

*wrong

OpenStudy (anonymous):

\[f ^{-1}(x)=\]

OpenStudy (anonymous):

the answer should look like this

OpenStudy (anonymous):

hey where'd u go? lol

OpenStudy (anonymous):

replace y with x

OpenStudy (anonymous):

that is correct! :D could you show me the steps please so i can write them down?

OpenStudy (anonymous):

let us say f(x)=y. y-5=sqrt(2x-5) (y-5)^2=2x-5 [squaring on both sides..] solve for x from here on..

OpenStudy (anonymous):

ok thanks. i need help with another one so if u can, please help. thanks again.